Carbonyl fluoride, COF2, is an important intermediate used in the production of fluorine-containing compounds. For instance, it is used to make the refrigerant carbon tetrafluoride, CF4 via the reaction

2COF2(g)⇌CO2(g)+CF4(g),    Kc=8.90

If only COF2 is present initially at a concentration of 2.00 M, what concentration of COF2 remains at equilibrium?

Express your answer with the appropriate units.
Answer #1

Given, the equilibrium reaction,

2COF2(g) \rightleftharpoons CO2(g) + CF4(g) ----------- Kc = 8.90

Initial concentration of COF2 = 2.00 M

Drawing an ICE chart,

2COF2 CO2(g) CF4(g)
I(M) 2.00 0 0
C(M) -2x +x +x
E(M) 2.00-2x x x

Kc = [CO2] [CF4] / [COF2]2

8.90 = [x] [x] / [2.00-2x]

8.90 = x2 / [2.00-2x]

17.8 - 17.8 x - x2 = 0

x2 + 17.8 x -17.8 = 0

Solving the quadratic equation,

x = 0.9494

We know,

[COF2]eq = 2.00-2x

[COF2]eq = 2.00-2(0.9494)

[COF2]eq = 0.101 M